OpenStudy (anonymous):

I have the answer can someone check me please... Hugh is driving a car down the road. Let d(t) give the distance he is from home t hours after he left. If d(t)=-1.5t^2+64t find Hugh's acceleration 4 hours after he left Acceleration is?

zepdrix (zepdrix):

Acceleration is the second derivative of displacement. What did you get for an answer farmgirl? c:

OpenStudy (anonymous):

I got 280

zepdrix (zepdrix):

So it seems you found Hugh's `distance` after 4 hours. \(\large d(4)=-1.5(4)^2+64(4)=232\) There was a small mistake in your calculation I think. Doesn't matter though, we don't want the distance, we want the acceleration at 4 hours. \[\large d'(t)=v(t) \qquad \rightarrow \qquad v(t)=-3t+64\]The second derivative will be \(\large d''(t)=a(t)\), the acceleration function that we're looking for. Think you can find the second derivative? After you do, plug in \(\large t=4\).

OpenStudy (anonymous):

so then is the answer 52

zepdrix (zepdrix):

No, it appears you plugged 4 into the velocity function. Take the derivative again to get the acceleration function, and then plug 4 into it.

OpenStudy (phi):

this is a calculus problem. Do you understand how to take the second derivative ?

OpenStudy (anonymous):

I am confussed

OpenStudy (phi):

can you take the derivative of d(t)= -1.5t^2+64t with respect to t ? zepdrix already did it for you , and got v(t) = -3t + 64 now you have to take the derivative of v(t) with respect to t can you do that ?

OpenStudy (anonymous):

-3t+64t

OpenStudy (phi):

ok that is the first derivative. now take the derivative of that

OpenStudy (anonymous):

how do you do that

OpenStudy (phi):

first it is -3t+64 (not -3t+64t) the derivative of -3t is ? and derivative of 64 is ?

OpenStudy (phi):

\[ \frac{d}{dt} (-3t) = ?\]

OpenStudy (anonymous):

-12 and 256

OpenStudy (phi):

do you know how to take the derivative of -3t where t is the variable ?

OpenStudy (phi):

do you know how to do \[ \frac{d}{dx} 3x \]

OpenStudy (phi):

If not, you really can't do this problem.

OpenStudy (anonymous):

wait just redid it -85.33

OpenStudy (anonymous):

then 1024

OpenStudy (anonymous):

Is that right

OpenStudy (phi):

I guess you don't understand. To do this problem you take the derivative of d(t)= -1.5t^2+64t you get v(t) = -3t + 64 (this says d distance/dt = velocity) the change in velocity with respect to time is acceleration a(t)= dv/dt = -3 that says the acceleration is always -3 (what ever units they are using)

OpenStudy (phi):

If that does not make sense, you have to brush up on your calculus (and learn that velocity is the first derivative of distance, and acceleration is the 2nd derivative of distance)

OpenStudy (anonymous):

ok thanks